OpenStudy (theedwardsfamily):

Which fractions are equivalent to -16/24 ? @Nnesha Choose all answers that are correct. A. -2/3 B. -5/8 C. -32/48 D. -6/9

OpenStudy (juicstice):

what do 16 and 24 have in common? or what are they divisible by?

OpenStudy (theedwardsfamily):

umm....4 and 8?

OpenStudy (juicstice):

well yes but that is not an answer so 8 is divisible by 16 so what x 8 equals 16 what x 8 equals 24

OpenStudy (dtan5457):

@TheEdwardsFamily do you get this?

OpenStudy (theedwardsfamily):

yes i guess

OpenStudy (dtan5457):

you need to see that \[-\frac{ 16 }{ 24 }\div \frac{ 8 }{ 8 }=-\frac{ 2 }{ 3 }\] and that \[-\frac{ 16 }{ 24 }\times \frac{ 2 }{ 2 }=-\frac{ 32 }{ 48 }\]

OpenStudy (dtan5457):

that are equal ^^

OpenStudy (theedwardsfamily):

2*8= 16 and 4*6=24

OpenStudy (dtan5457):

what are you referring to..?

OpenStudy (theedwardsfamily):

what 16 and 24 are divisble by @dtan5457

OpenStudy (juicstice):

well you did get the first equation right but what x 8 = 24 because if you want to find what the fraction is equivalent to you have to do the same to the top and the bottom so divide them both by 8 and you will get one answer

OpenStudy (dtan5457):

16 is not divisible by 6

OpenStudy (theedwardsfamily):

3*8

OpenStudy (juicstice):

yes. so one of the answers is -2/3 because 16 and 24 are both divisible by 8 and it has to be negative because the original fraction is negative

OpenStudy (juicstice):

I will tell you there are 2 more answers that are possible

OpenStudy (theedwardsfamily):

I gotta get ready for dinner

OpenStudy (theedwardsfamily):

can we come back 2mrw around 10am?

OpenStudy (juicstice):

I don't think I can

OpenStudy (juicstice):

I can just give you the answers real fast if you need to go

OpenStudy (juicstice):

it's A, C, and D

OpenStudy (theedwardsfamily):

thanks